# Structured Initialization for Vision Transformers

### v1

@@ 5 Experiments and Analysis
In this section, we show that our impulse initialization can improve the performance of ViT on small-scale datasets in Sec. 5.1, while it does not limit the learning flexibility of ViT models on large-scale datasets in Sec. 5.2. In Sec. 5.5, we show that even for a pretrained method like weight selection [33], the pretrained model initialized with our impulse structure can provide better pretrained weights with a relatively smaller-scale dataset. Finally, in Sec. 5.3, we show that besides ViTs, the concept of introducing architectural bias into initializations can also be generalized to other models like Swin Transformers and MLP-Mixer. Note that all experiments were conducted on a single node with 8 Tesla V100 SXM3 GPUs, each with 32GB of memory. Specifically, all the experiments on the small-scale datasets took about three hours to train each model, while the experiments on the ImageNet-1K took about two days. Please note that all results were reported based on our experiments with retrained models. As a result, minor discrepancies may exist between our reported results and those in published papers. However, the key focus of this analysis remains on the improvements achieved through different initialization strategies.
@@ 5.1 Small and Medium-Scale Datasets
[…]
The experimental results are presented in Tab. 1. Our method consistently yields substantial improvements over the default initialization across all evaluated datasets. On medium-scale datasets, it achieves performance gains of $2\%{\sim}5\%$, while on small-scale datasets, improvements can reach approximately $8\%$, and in some cases, up to $20\%$. While mimetic initialization also improves the performance, our impulse initialization shows superior efficacy, attributed to the convolutional structure integrated in the attention initialization. Notably, as the dataset size decreases, the performance gap between our method and the default (or mimetic) initialization gets larger. This observation validates that the convolutional inductive bias introduced by our initialization becomes increasingly important when there is less data for the model to learn the spatial dependencies.
@@
5.2 Large-Scale Datasets
In this section, we compare the performance of ViT-Tiny, ViT-Small, and ViT-base with default, mimetic, and our impulse initialization on a large-scale dataset—ImageNet-1K (over 1M training images). We follow the training recipe from DeiT [26]—a classic and efficient training recipe for training ViT models on ImageNet-1K. We directly use the ViT structure and training codes in the timm library, except for adding our implementation of initialization. All the experiments were trained with the same hyperparameters.
We show the comparison results on the ImageNet-1K dataset in Tab. 2. Detailed training hyperparameters and training curves can be found in the appendix. Our method maintains comparable performance with default initialization, demonstrating that the convolutional inductive bias introduced during initialization does not hinder the model’s flexibility in learning data-driven dependencies. This indicates that while the transformer architecture begins training with structurally imposed spatial priors, the attention mechanism retains full capacity to learn optimal feature representations when sufficient training data is available. Furthermore, the convolutional structure we introduced in the attention map initialization not only accelerates early convergence but also improves the robustness to variations in training hyperparameters. We will provide more quantitative results in the appendix.
@@ 5.2 Large-Scale Datasets
indicates large-scale datasets.
Method Model $\blacksquare$  ViT-Tiny $\blacksquare$  ViT-Small $\blacksquare$  ViT-Base   Default [31] 72.71   2.27$\downarrow$ 79.68   2.27$\downarrow$ 76.25111Although we strictly follow the training recipe reported in DeiT, we acknowledge that our baseline performance appears slightly lower. This discrepancy may stem from subtle differences in training code or hardware configurations. However, the primary purpose of this table is to demonstrate that our initialization does not compromise the performance of large models.  2.27$\downarrow$   Mimetic [28] 72.90   0.19$\uparrow$ 80.26   0.58$\uparrow$ 79.12   2.87$\uparrow$   Ours (impulse) 72.76   0.05$\uparrow$ 80.40   0.72$\uparrow$ 80.48   4.23$\uparrow$
@@
Method Model&Data Swin-B on $\blacksquare$  ImageNet-1K Swin-T on $\blacktriangle$ CIFAR-10 MLP-Mixer on $\blacktriangle$ CIFAR-10 Default [31] 83.14   2.27$\downarrow$ 89.85   2.27$\downarrow$ 87.00   2.27$\downarrow$ Mimetic [28] 83.14      89.24   0.61$\downarrow$ Ours (impulse) 83.55   0.41$\uparrow$ 91.19   1.34$\uparrow$ 88.78   1.78$\uparrow$
@@ 5.3 ViT Variants
We evaluated our approach with Swin Transformer-Base (Swin-B) architecture on the ImageNet-1K following the training recipe in the original Swin Transformer paper [17]. To further demonstrate the flexibility of our approach, we applied our impulse initialization to the MLP-Mixer. For experiments with Swin Transformer-Tiny (Swin-T) and MLP-Mixer, we follow training settings in [34], which is specifically designed for training different models on CIFAR-10.
The results are summarized in Tab. 3. For models with strong learning capacity but limited inherent inductive biases at initialization, introducing a structured initialization consistently enhances performance without compromising the model’s capacity to learn complex data dependencies. In particular, while Swin Transformer incorporates convolutional inductive bias through windowed self-attention, applying our structured initialization further improves the performance by $1.34\%$ on smaller-scale datasets such as CIFAR-10, with no degradation on large-scale datasets like ImageNet-1K. MLP-Mixer, which replaces the spatial convolutions in ConvMixer with MLP layers, typically struggles to train effectively on small-scale datasets. However, initializing the spatial MLPs with a convolutional structure leads to a $1.78\%$ performance gain on CIFAR-10. In contrast, mimetic initialization—designed to replicate empirical weight distributions from pretrained ViTs—shows negligible benefits or degrades performance, highlighting its limited generalizability outside the specific pretrained ViT structures.
@@ 5.4 Attention Maps
The initialized attention maps produced by default, mimetic, and our impulse initializations are shown in Fig. 3. For better visualization, we use $32{\times}32$ CIFAR-10 images as input with a $4{\times}4$ patch size, yielding the token size as $8{\times}8{=}64$. We observe that the default initialization generates near-identical attention values with little distinguishable spatial structure, while the mimetic initialization introduces diagonally dominant values by adding an identity matrix during initialization. Notably, these two initialization methods initialize all the attention heads in the same way. In contrast, our initialization assigns different impulse patterns to each attention head, producing spatially diverse activations with off-diagonal peaks on specific impulse positions.
@@ 5.5 Pre-trained Model
In this section, we demonstrate that our initialization method remains compatible with existing pretraining pipelines. Building on the work of Xu et al. [33], who showed that weight selection from ImageNet-21K-pretrained ViT-Small models effectively initializes ViT-Tiny architectures, we extend this approach to ImageNet-1K pertaining. We pretrained three ViT-Small models on ImageNet-1K using three initialization strategies: default, mimetic, and our proposed impulse initialization. From these, we derived three ViT-Tiny initialization variants—termed “1K-default”, “1K-mimetic”, and “1K-impulse”—using the same weight selection methodology on smaller-scale datasets. When training these initialized ViT-Tiny models on small and medium-scale datasets (results shown in Tab. 4), we observed that: (1) Switching from ImageNet-21K to ImageNet-1K—less training data—pretraining with default initialization typically incurs a performance drop of ${\sim}1\%$. (2) Mimetic initialization fails to mitigate this data degradation, (3) The ImageNet-1K pretraining model with our impulse-based initialization achieves performance parity with ImageNet-21K pretrained baselines. This highlights the robustness of our method even under reduced pretraining data regimes.
Table 4: Classification accuracy of ViT-Tiny on small-scale datasets with the weight selection method. Here, the weights selected for the experiments were pretrained on the ImageNet-1K (shortened as 1K, shown above the purple dashed line) and ImageNet-21K (shortened as 21K, shown below the purple dashed line) datasets with different initialization methods. Please note that pretraining on ImageNet-21K with default initialization is the original weight selection method [33]. We compare the performance to the default initialization method pretrained on ImageNet-1K (shaded in gray).
@@ 6 Conclusion
Supplemental Material: Structured Initialization for Vision Transformers
Appendix A Limitations and Broader Impacts
There exist several limitations in our initialization method: (1) Positional encoding.   Although positional encoding is a natural and effective choice for pseudo-inputs—due to its simplicity and data-independent nature—even simpler alternatives may exist for initializing the $\mathbf{Q}$ and $\mathbf{K}$ matrices. In this work, we focus solely on the initialization of $\mathbf{Q}$ and $\mathbf{K}$, but a more comprehensive initialization strategy that also considers patch embeddings and positional encodings could offer greater control over the structure of attention maps. Notably, since positional encoding is only added at the input layer, the influence on attention maps may diminish with increasing network depth, as illustrated in Fig. 3. (2) Hard constraints. Our Corollary 2 of Proposition 1 is based on the presumption that the filters are $f^{2}$$k$ spanned, which is usually a characteristic inherent in CNNs. However, in ViTs, the limited number of heads may be inadequate to span the filter space of a small kernel. Finding better adaptations in this scenario remains a challenge. (3) Value initialization. Our method does not consider the initialization for the value weights $\mathbf{V}$ and the projection matrix.
Appendix B Frequently Asked Questions
@@
A: Our method focuses on introducing the architectural inductive bias from CNNs to initialize the attention map without changing the Transformer architectures. We have emphasized this argument in the abstract and introduction, stating that unlike previous arts [35, 15, 7, 32, 16] that directly introduce convolutions into attention, potentially damaging the structural advantages of Transformers, our method only introduces architectural inductive bias in attention map initialization, maintains the inherent structural flexibility in ViTs. Therefore, our method preserves the Transformer architectures and still allows the ViTs to learn flexible, dynamic global relationships. This is also one of the central innovations of our method. In addition, we have also provided experiments on large-scale applications to validate the stable performance of our method that preserves the architectural flexibility of ViTs.
@@
A: As stated in Sec. 1, we would like to emphasize the advantages of our proposed structured initialization method: 1) Unlike previous methods that do transfer learning on large-scale pre-trained ViTs, our method involves no pre-training of large-scale models on large-scale datasets, which may not be readily available; 2) Our method shares the advantages of both CNNs and ViTs.
In general, introducing inductive bias in CNNs for initializing the attention map helps ViTs begin learning from a more reasonable/stable starting point, while the default random initialization can lead to a noisy starting point, especially when training on small-scale datasets. Notably, our initialization method does not alter the ViT structure, maintaining the advantages of the Transformer architectures in learning dynamic, long-dependent global features. In contrast, methods that directly combine the architectures of CNNs and ViTs alter the Transformer architectures, potentially compromising its architectural advantages. Please also refer to the theoretical analysis of random filters in Sec. 3 and the convolutional representation matrix in Appendix D of the main paper for more theoretical explanations.
For pretrained models, we noted in the introduction section that their reliance on access to pretrained models makes them impractical for domain-specific scenarios. For example, pretraining a large model on extensive datasets only to deploy a smaller model on limited data is inefficient and often unreasonable. Furthermore, their effectiveness heavily depends on the performance of the specific model architectures, offering no inherent advantages from using pretrained weights. Nevertheless, even on pretrained models, our initialization still outperforms other methods. For additional quantitative evidence, please refer to the experimental analysis in Sec. 5.1 and Sec. 5.5.
@@
A: While one might expect that a rigid initialization could limit the model’s flexibility, our experimental results in Sec. 5 demonstrate that the impulse structure does not hurt the training. On the contrary, it consistently improves the performance, particularly on small and medium-scale datasets. This is attributed to the structured initialization introducing a beneficial inductive bias, which guides the model toward learning useful representations in the early stage of training.
@@
Appendix C Proof for Proposition 1

### v2

@@ 5 Experiments and Analysis
In this section, we show that our impulse initialization can improve the performance of ViT on small-scale datasets in Sec.˜5.1, while it does not limit the learning flexibility of ViT models on large-scale datasets in Sec.˜5.2. In Sec.˜5.6, we show that even for a pretrained method like weight selection [xu2024initializing], the pretrained model initialized with our impulse structure can provide better pretrained weights with a relatively smaller-scale dataset. Finally, in Sec.˜5.4, we show that besides ViTs, the concept of introducing architectural bias into initializations can also be generalized to other models like Swin Transformers and MLP-Mixer. Note that all experiments were conducted on a single node with 8 Tesla V100 SXM3 GPUs, each with 32GB of memory, if not specified. Specifically, all the experiments on the small-scale datasets took about three hours to train each model, while the experiments on the ImageNet-1K took about two days. Please note that all results were reported based on our experiments with retrained models. As a result, minor discrepancies may exist between our reported results and those in published papers. However, the key focus of this analysis remains on the improvements achieved through different initialization strategies.
@@ 5.1 Small and Medium-Scale Datasets
In this section, we compare the performance of ViT-Tiny under three initialization strategies: the default [rw2019timm], mimetic [trockman2023mimetic], and our impulse initialization. Experiments are conducted on medium-scale datasets ($\sim$}50K training images), including Food-101 [bossard2014food], CIFAR-10, and CIFAR-100 [krizhevsky2009learning], as well as small-scale datasets ($\sim$}5K training images), such as STL-10 [coates2011analysis], Flowers [Nilsback08], and Pets [parkhi12a]. We follow the training recipe from [xu2024initializing], which is proven to be useful in training ViT-Tiny on small and medium-scale datasets. Their codes are based on the timm library [rw2019timm]. By default, all the weights are initialized with a truncated normal distribution. For a fair comparison, all the experiments were run with identical codes except for the choice of initialization methods. Please also note that although [xu2024initializing] initializes the model weights from large pretrained models, we did not adopt any pretraining step for this experiment. In contrast, we apply default, mimetic, and our impulse initialization methods to ViT-Tiny models and training these models from scratch.
The experimental results are presented in Tab.˜1. For simplicity, we keep the statistical results on using different seeds with stochastic filter generations with 5 different runs in Appendix˜F and Fig.˜6. Our method consistently yields substantial improvements over the default initialization across all evaluated datasets. On medium-scale datasets, it achieves performance gains of $2\%{\sim}5\%$, while on small-scale datasets, improvements can reach approximately $8\%$, and in some cases, up to $20\%$. While mimetic initialization also improves the performance, our impulse initialization shows superior efficacy, attributed to the convolutional structure integrated in the attention initialization. Notably, as the dataset size decreases, the performance gap between our method and the default (or mimetic) initialization gets larger. This observation validates that the convolutional inductive bias introduced by our initialization becomes increasingly important when there is less data for the model to learn the spatial dependencies.
@@
indicates large-scale datasets. Please note that for the last column (shaded in yellow), we report the experimental results on ViT-Base with specific settings to make the default initialization-based ViT comparable to the results reported in concurrent papers.
[…]
5.2 Large-Scale Datasets
In this section, we compare the performance of ViT-Tiny, ViT-Small, and ViT-base with default, mimetic, and our impulse initialization on a large-scale dataset—ImageNet-1K (over 1M training images). We follow the training recipe from DeiT [touvron2021training]—a classic and efficient training recipe for training ViT models on ImageNet-1K. We directly use the original ViT structure and training codes in the timm library, except for adding our implementation of initialization. All the models were trained with the same hyperparameters starting from scratch without any pretraining or distillation.
We show the comparison results on the ImageNet-1K dataset in Tab.˜2. Detailed training hyperparameters and training curves can be found in Appendix˜E. In particular, we find that the rapid update for the baseline ViT codebase (i.e., timm library) and the difference in GPU hardware settings result in a small discrepancy in the default initialization-based ViT-Base model between our main result (${\sim}81.24$) and the results reported in concurrent papers (${\sim}81.89$). Therefore, we have included an additional column (shaded in yellow) in Tab.˜2 for ViT-Base model that uses 16 Tesla V100 GPUs with an extra 0.3 color jittering data augmentation for a clearer comparison.
Despite different training settings and comparisons, our method maintains comparable performance with default initialization, demonstrating that the convolutional inductive bias introduced during initialization does not hinder the model’s flexibility in learning data-driven dependencies. This indicates that while the transformer architecture begins training with structurally imposed spatial priors, the attention mechanism retains full capacity to learn optimal feature representations when sufficient training data is available. Furthermore, the convolutional structure we introduced in the attention map initialization not only accelerates early convergence but also improves the robustness to variations in training hyperparameters. We provide more quantitative results in Appendix˜E.
Figure 3: Training curves of ViT-Tiny on CIFAR-10 and ViT-Base on ImageNet-1K using default, mimetic, and impulse initialization. The zoomed-in box shows the training curve in the final training stage from epoch 200 to epoch 300.
5.3 Training Curves and Analysis
In Fig.˜3, we show the training accuracy curves across $300$ epochs for different initialization methods of ViT-Tiny on CIFAR-10 and ViT-Base on ImageNet-1K. For the small model on medium-scale dataset, our impulse initialization consistently outperforms the default or mimetic initialization throughout the entire training process. On the large-scale dataset with training large ViT-Base model, our impulse initialization method and the mimetic initialization have shown faster convergence rate than the default initialization at the beginning of the training. However, the mimetic initialization shows degraded performance even to the default initialization at the last $100$ training epochs, indicating limited ability in large-scale model training. On the contrary, our method does not limit the learning ability of the large-scale ViT models, showing a significant advantage in the final training stage where the performance surpass both the default and the mimetic initialization methods.
@@ 5.3 Training Curves and Analysis
Figure 4: Visualization of attention maps in ViT-T using ours, mimetic [trockman2023mimetic], and default [xu2024initializing] initializations. Red boxes highlight zoomed-in details of the 16× $16\,{\times}\,$16 upper left corner in attention maps. White boxes indicate the 8× $8\,{\times}\,$8 sub-blocks of the zoomed-in attention maps. Our structured initialization method offers distinct attention peaks aligned with the impulse structures across different heads. Head 1 offers a peak at +1+1 offset from the main diagonal. Head 2 offers a peek at 8-8 offset (equivalent to $\text{image\_size}$) from the main diagonal. Both mimetic and random initialization methods initialize all the attention heads identically. Specifically, mimetic initialization primarily strengthens the main diagonal of the attention map for each head, while random initialization shows minimal structural patterns with flatter peak values.
@@ 5.4 ViT Variants
We evaluated our approach with Swin Transformer-Base (Swin-B) architecture on the ImageNet-1K following the training recipe in the original Swin Transformer paper [liu2021swin]. To further demonstrate the flexibility of our approach, we applied our impulse initialization to the MLP-Mixer. For experiments with Swin Transformer-Tiny (Swin-T) and MLP-Mixer, we follow training settings in [yoshioka2024visiontransformers], which is specifically designed for training different models on CIFAR-10.
The results are summarized in Tab.˜3. For models with strong learning capacity but limited inherent inductive biases at initialization, introducing a structured initialization consistently enhances performance without compromising the model’s capacity to learn complex data dependencies. In particular, while Swin Transformer incorporates convolutional inductive bias through windowed self-attention, applying our structured initialization further improves the performance by $1.34\%$ on smaller-scale datasets such as CIFAR-10, with no degradation on large-scale datasets like ImageNet-1K. MLP-Mixer, which replaces the spatial convolutions in ConvMixer with MLP layers, typically struggles to train effectively on small-scale datasets. However, initializing the spatial MLPs with a convolutional structure leads to a $1.78\%$ performance gain on CIFAR-10. In contrast, mimetic initialization—designed to replicate empirical weight distributions from pretrained ViTs—shows negligible benefits or degrades performance, highlighting its limited generalizability outside the specific pretrained ViT structures.
@@ 5.5 Attention Maps
The initialized attention maps produced by default, mimetic, and our impulse initializations are shown in Fig.˜4. For better visualization, we use $32{\times}32$ CIFAR-10 images as input with a $4{\times}4$ patch size, yielding the token size as $8{\times}8{=}64$. We observe that the default initialization generates near-identical attention values with little distinguishable spatial structure, while the mimetic initialization introduces diagonally dominant values by adding an identity matrix during initialization. Notably, these two initialization methods initialize all the attention heads in the same way. In contrast, our initialization assigns different impulse patterns to each attention head, producing spatially diverse activations with off-diagonal peaks on specific impulse positions.
@@ 5.6 Pre-trained Model
In this section, we demonstrate that our initialization method remains compatible with existing pretraining pipelines. Building on the work of Xu et al[xu2024initializing], who showed that weight selection from ImageNet-21K-pretrained ViT-Small models effectively initializes ViT-Tiny architectures, we extend this approach to ImageNet-1K pertaining. We pretrained three ViT-Small models on ImageNet-1K using three initialization strategies: default, mimetic, and our proposed impulse initialization. From these, we derived three ViT-Tiny initialization variants—termed “1K-default”, “1K-mimetic”, and “1K-impulse”—using the same weight selection methodology on smaller-scale datasets. When training these initialized ViT-Tiny models on small and medium-scale datasets (results shown in Tab.˜4), we observed that: (1) Switching from ImageNet-21K to ImageNet-1K—less training data—pretraining with default initialization typically incurs a performance drop of ${\sim}1\%$. (2) Mimetic initialization fails to mitigate this data degradation, (3) The ImageNet-1K pretraining model with our impulse-based initialization achieves performance parity with ImageNet-21K pretrained baselines. This highlights the robustness of our method even under reduced pretraining data regimes.
Table 4: Classification accuracy of ViT-Tiny on small-scale datasets with the weight selection method. Here, the weights selected for the experiments were pretrained on the ImageNet-1K (shortened as 1K, shown above the purple dashed line) and ImageNet-21K (shortened as 21K, shown below the purple dashed line) datasets with different initialization methods. Please note that pretraining on ImageNet-21K with default initialization is the original weight selection method [xu2024initializing]. We compare the performance to the default initialization method pretrained on ImageNet-1K (shaded in gray).
@@
6 Limitations and Broader Impacts
There exist several limitations in our initialization method: (1) Positional encoding.   Although positional encoding is a natural and effective choice for pseudo-inputs—due to its simplicity and data-independent nature—even simpler alternatives may exist for initializing the $\mathbf{Q}$ and $\mathbf{K}$ matrices. In this work, we focus solely on the initialization of $\mathbf{Q}$ and $\mathbf{K}$, but a more comprehensive initialization strategy that also considers patch embeddings and positional encodings could offer greater control over the structure of attention maps. Notably, since positional encoding is only added at the input layer, the influence on attention maps may diminish with increasing network depth, as illustrated in Fig.˜4. (2) Hard constraints. Our Corollary 2 of Proposition 1 is based on the presumption that the filters are $f^{2}$$k$ spanned, which is usually a characteristic inherent in CNNs. However, in ViTs, the limited number of heads may be inadequate to span the filter space of a small kernel. Finding better adaptations in this scenario remains a challenge. (3) Value initialization. Our method does not consider the initialization for the value weights $\mathbf{V}$ and the projection matrix.
@@ 7 Conclusion
Appendix: Structured Initialization for Vision Transformers
Appendix A Frequently Asked Questions
@@
A: Our method focuses on introducing the architectural inductive bias from CNNs to initialize the attention map without changing the Transformer architectures. We have emphasized this argument in the abstract and introduction, stating that unlike previous arts [yuan2021incorporating, li2023uniformer, dai2021coatnet, wu2021cvt, liu2021efficient] that directly introduce convolutions into attention, potentially damaging the structural advantages of Transformers, our method only introduces architectural inductive bias in attention map initialization, maintains the inherent structural flexibility in ViTs. Therefore, our method preserves the Transformer architectures and still allows the ViTs to learn flexible, dynamic global relationships. This is also one of the central innovations of our method. In addition, we have also provided experiments on large-scale applications to validate the stable performance of our method that preserves the architectural flexibility of ViTs.
@@
A: As stated in Sec.˜1, we would like to emphasize the advantages of our proposed structured initialization method: 1) Unlike previous methods that do transfer learning on large-scale pre-trained ViTs, our method involves no pre-training of large-scale models on large-scale datasets, which may not be readily available; 2) Our method shares the advantages of both CNNs and ViTs.
In general, introducing inductive bias in CNNs for initializing the attention map helps ViTs begin learning from a more reasonable/stable starting point, while the default random initialization can lead to a noisy starting point, especially when training on small-scale datasets. Notably, our initialization method does not alter the ViT structure, maintaining the advantages of the Transformer architectures in learning dynamic, long-dependent global features. In contrast, methods that directly combine the architectures of CNNs and ViTs alter the Transformer architectures, potentially compromising its architectural advantages. Please also refer to the theoretical analysis of random filters in Sec.˜3 and the convolutional representation matrix in Appendix˜C of the main paper for more theoretical explanations.
For pretrained models, we noted in the introduction section that their reliance on access to pretrained models makes them impractical for domain-specific scenarios. For example, pretraining a large model on extensive datasets only to deploy a smaller model on limited data is inefficient and often unreasonable. Furthermore, their effectiveness heavily depends on the performance of the specific model architectures, offering no inherent advantages from using pretrained weights. Nevertheless, even on pretrained models, our initialization still outperforms other methods. For additional quantitative evidence, please refer to the experimental analysis in Sec.˜5.1 and Sec.˜5.6.
@@
A: While one might expect that a rigid initialization could limit the model’s flexibility, our experimental results in Sec.˜5 demonstrate that the impulse structure does not hurt the training. On the contrary, it consistently improves the performance, particularly on small and medium-scale datasets. This is attributed to the structured initialization introducing a beneficial inductive bias, which guides the model toward learning useful representations in the early stage of training.
@@
Appendix B Proof for Proposition 1